Two towns A and B are 220 km apart. A bus left town A at 11.00 a.m. and travelled towards B at an average speed of 60 km/h. At the same time, a matatu left town B for A at an average speed of 80 km/h. The matatu stopped for a total of 45 minutes on the way before meeting the bus. Calculate the distance covered by the bus before meeting the matatu.
Answer by greenestamps(13200)



During the 45 minutes (3/4 hour) that the matatu is stopped, the bus travels (3/4)(60) = 45km. So the total distance the two vehicles travel while both are moving is 220-45 = 175km.

The rate at which they approach each other is the sum of their two rates, which is 60+80 = 140km/hr.

The time required to cover the 175km at a rate of 140km/h is 175/140 = 5/4 hours.

So when the bus meets the matatu it has been traveling for 3/4+5/4 = 8/4 = 2 hours; the distance it has traveled at that time is 2(60) = 120km.